Write the equations in logarithmic form.
(a) 729 =(^3
(b) 64= (1/8) ^-2
(c) a=B^c
1
Expert's answer
2013-02-20T09:26:03-0500
(a) 729 =9^3
In this example, the base is 4 and the base moved from the
right side of the exponential equation to the left side of the
logarithmic equation and the word “log” was added:
log 9 729 = 3

(b) 64= (1/8) ^-2
the base is 1/8:
log 1/8 64 = -2

(c) a=B^c
the base is B:
log B a = c